1. Pencil “lead” is almost pure graphite.  Graphite is the stable elemental form of carbon at 25°C and 1 atm.  Diamond is an allotrope of graphite. Given diamond: ΔHf° = 1.9 kJ/mol; S° = 2.4 J/mol*K, at what temperature are the two forms in equilibrium at 1 atm?

C (graphite) C (diamond)


1
Expert's answer
2021-01-14T04:25:17-0500

At . Equilibrium

ΔG° = 0 .

ΔGo = ΔHf° - T . ΔS° .

0 = 1.9 × 1000 - T × 2.4

So, T = 1900 / 2.4 = 4560 K .